OpenStudy (anonymous):

Why is 5 ⅓ not equal to 5.3?

OpenStudy (radar):

1/3 does not = 3/10

OpenStudy (anonymous):

because the fraction 1/3 has a repeating decimal whereas .3 is the equivalent of 3/10 and does not continue on forever.

OpenStudy (radar):

5 1/3 =16/3 5 3/10 =53/10 16/3 does not equal 53/10